In a physics lab students are conducting an experiment to learn about the heat capacity of different materials. The first group is instructed to add 1.5-g lead pellets at a temperature of 92°C to 345 g of water at 16°C. A second group is given the same number of 1.5-g pellets as the first group, but these are now aluminum pellets. Assume that no heat is lost to or gained from the surroundings for either group. (a) If the final equilibrium temperature of the lead pellets and water is 25°C, how many whole pellets did the first group use in the experiment? The specific heat of lead is 0.0305 kcal/(kg · °C). ___ pellets (b) Will the final equilibrium temperature for the second group be higher, lower, or the same as for the first group? The specific heat of aluminum is 0.215 kcal/(kg · °C). -higher -lower or - the same (c) What is the equilibrium temperature of the aluminum and water mixture for the second group? (°C)
Answer #1

+ Olesed (T: -T6 ) = mw Cw (Tf - Tiw) mw C (Tf -Tiw) mc dead Ti-Tf) 345(1)(25-16°c) 1.5 (0.005) (92-25) = 1013 16 hisher Appl

1.5 (0.005) 92-28) - 1013 (6) hisher Apply principle of Calormelry -Qsain MA - MAICA (92-7)= mw (w (7-14 T = MACA (92) +mw w